# TOMATOES: Topology and Material Optimization for Latent Heat Thermal Energy Storage Devices

### v1

@@ 4.4 Variation in Discharge Time
Next, we consider the allowed discharge time for the design of the LHTES systems. We consider four scenarios with different discharge times of 55, 110, 165 and 220 hours.
The results are illustrated in Figure˜13. As expected, the total energy discharged increases as the allowed discharge time increases. Furthermore, we observe that the geometry of the HCM fins are thicker and concentrated near the heat source for shorter durations. As the duration extends, the fins branch out more deeply into the PCM domain, creating finer, more complex networks. This allows for more effective heat extraction from regions farther from the source over the longer periods.
@@
Interestingly, the framework also changes the material selection based on the time allowed. For shorter durations ($t=55,110$ hours), it selects Graphite as the HCM. For longer durations ($t=165,220$ hours), it switches to Magnesium. This suggests that different materials are optimal for different heat discharge rates and durations.
@@ 5 Conclusion
This work introduces a computational framework for the co-optimization of geometry and material of latent heat thermal energy storage (LHTES) devices. In particular, we utilize a variational autoencoder (VAE) to represent the discrete material databases of high-conductivity materials (HCMs) and phase-change materials (PCMs) as continuous, differentiable latent spaces ([7]). Then, we integrate these differentiable material representations for the simultaneous optimization of material choice and geometric layout. We performed several numerical experiments to validate our central hypothesis: the co-design yields LHTES systems with superior performance compared to conventional sequential design workflows.
The present study offers several avenues for future research. First, the thermal analysis included only heat conduction, neglecting the effects of convection within the molten PCM. Second, the material databases used in this work were primarily for illustration. The practical implementation of this framework will require the use of comprehensive, experimentally validated material datasets that account for property uncertainties and temperature dependencies. Third, the optimization formulation did not consider constraints such as manufacturability, material compatibility, or cyclic stability, which are essential for real-world deployment. Fourth, we observed that the optimizer can become trapped in local solutions, often requiring multiple initializations. One direction for future work can be to improve global-search robustness and reduce sensitivity to initialization to better approach optimal designs. Fifth, we optimized the LHTES in a 2D domain and focused solely on the selection of the HCM and PCM. As real systems use a circulating HTF, a natural extension can be selecting the optimal HTF and extend the framework to 3D configurations. Finally, experimental validation of our optimized designs is also a crucial area for future investigation.

### v2

@@ 4.5 Variation in Discharge Time
Next, we consider the allowed discharge time for the design of the LHTES systems. We consider four scenarios with different discharge times of 5.5, 11, 16.5 and 22 hours.
The results are illustrated in Figure˜15. As expected, the total energy discharged increases as the allowed discharge time increases. Furthermore, we observe that the geometry of the HCM fins are thicker and concentrated near the heat source for shorter durations. As the duration extends, the fins branch out more deeply into the PCM domain, creating finer, more complex networks. This allows for more effective heat extraction from regions farther from the source over the longer periods.
@@
Interestingly, the framework also changes the material selection based on the time allowed. For shorter durations ($t=5.5,11$ hours), it selects Graphite as the HCM. For longer durations ($t=16.5,22$ hours), it switches to Magnesium. This suggests that different materials are optimal for different heat discharge rates and durations.
@@ 5 Conclusion
This work introduces a computational framework for the co-optimization of geometry and material of latent heat thermal energy storage (LHTES) devices. In particular, we utilize a variational autoencoder (VAE) to represent the discrete material databases of high-conductivity materials (HCMs) and phase-change materials (PCMs) as continuous, differentiable latent spaces ([chandrasekhar2022integrating]). Then, we integrate these differentiable material representations for the simultaneous optimization of material choice and geometric layout. We performed several numerical experiments to validate our central hypothesis: the co-design yields LHTES systems with superior performance compared to conventional sequential design workflows.
The present study offers several avenues for future research. First, the thermal analysis included only heat conduction, neglecting the effects of convection within the molten PCM. Recent topology optimization studies incorporating natural convection (e.g., [wang2023topology]) have demonstrated that buoyancy forces can influence optimal designs. Specifically, for long charging durations where convection dominates, the optimizer favors simpler V-shaped geometries over complex branching. This geometrical shift occurs to minimize the volume of the HCM (which generally possesses lower specific heat capacity than the PCM), thereby maximizing the total energy storage capacity once convective mixing becomes effective. Extending our differentiable solver to include fluid dynamics is therefore a critical step toward capturing these higher-fidelity trade-offs. Second, the material databases used in this work were primarily for illustration. The practical implementation of this framework will require the use of comprehensive, experimentally validated material datasets that account for property uncertainties and temperature dependencies. Third, the optimization formulation did not consider constraints such as manufacturability, material compatibility, or cyclic stability, which are essential for real-world deployment. These factors would significantly influence the optimization result. For instance, specific material combinations (e.g., copper and fatty acids) can exhibit chemical degradation or corrosive interaction [ostry2020compatibility], while galvanic coupling between dissimilar metals (e.g., aluminum and copper) in salt hydrates can lead to structural failure [farrell2006corrosive]. Incorporating compatibility constraints would enable the optimizer to avoid these chemically unstable pairs. Similarly, incorporating constraints on cyclic degradation driven by volumetric expansion [gao2013prediction] and manufacturing limitations [liu2018current] would penalize the complex, thin-walled branching features, thereby favoring robust, printable geometries. Future exploration of this framework will address these practical constraints by incorporating compatibility masks into the latent space and by directly incorporating geometric length-scale constraints into the loss function. Fourth, we observed that the optimizer can become trapped in local solutions, often requiring multiple initializations. One direction for future work can be to improve global-search robustness and reduce sensitivity to initialization to better approach optimal designs. Fifth, we optimized the LHTES in a 2D domain and focused solely on the selection of the HCM and PCM. As real systems use a circulating HTF, a natural extension can be selecting the optimal HTF and extend the framework to 3D configurations. Sixth, while this study focused on maximizing energy discharge, the underlying differentiable framework is objective-agnostic. Future studies can leverage this flexibility to optimize for alternative objectives, such as minimizing spatial temperature gradients or targeting specific discharge rates, without modifying the core framework. Finally, future work will target the experimental validation of these designs using Laser Powder Bed Fusion (LPBF) additive manufacturing. As demonstrated recently [morciano2024enhanced], LPBF is capable of manufacturing the complex structures predicted by our co-design framework.
Appendices
Appendix A Numerical Validation
The numerical accuracy of the computational framework is derived from the finite element solver’s implementation of the transient, nonlinear heat equation with phase change. To verify accuracy, we performed a validation study against the classical one-dimensional Stefan problem. This benchmark serves to confirm the solver’s ability to capture the moving boundary position and the transient temperature distribution within the PCM.
The Stefan problem consists of a semi-infinite slab of PCM, initially solid at its melting temperature $T_{m}$. At time $t=0$, the boundary at $x=0$ is raised to a constant wall temperature $T_{w}>T_{m}$, initiating the melting process. Heat transfer is driven solely by conduction in the liquid phase. The problem has an analytical solution for the temperature distribution $T(x,t)$ and the melting front location $\delta(t)$ [ogoh2010stefan].
[…]
Figure˜16 compares the numerical temperature profiles against the analytical solution at various time instances ($t=1,3,6,12$ hours). The numerical results obtained with a mushy-zone width of $\Delta T=4$ K show good agreement with the analytical solution.
To ensure numerical convergence, we conducted grid independence and temporal sensitivity studies. The grid study evaluated mesh resolutions of $60\times 60$, $80\times 80$, and $100\times 100$ elements while maintaining a fixed time step of $\Delta t=60$ s (Figure˜17). Conversely, temporal sensitivity was assessed by testing time steps of $30$ s, $60$ s, and $90$ s with a fixed mesh resolution of $80\times 80$ elements (Figure˜16). Similar convergence studies were conducted to determine the appropriate grid resolution and time step size for the main optimization problem.
Figure 16: Temporal sensitivity study comparing temperature profiles for time step sizes of $30$ s, $60$ s, and $90$ s with a mesh resolution of $80\times 80$ elements.
Figure 17: Grid independence study comparing temperature profiles for mesh resolutions of $60\times 60$, $80\times 80$, and $100\times 100$ elements with a fixed time step of $\Delta t=60$ s.